Please help find the missing factor...
4/x+2= ?/(x+2)(x+1)
What is missing?
I appreciate your help.

Answer by dabanfield(803)

4/x+2= ?/(x+2)(x+1)
To solve for ? we can eliminate the denominator on the right by multiplying both sides by (x+2)*(x+1):
Then we have:
[4/(x+2)] = ?
4*(x+1) =
4*1*(x+1) =
4*(x+1) = ?